# Bit Threads: From Entanglement to Geometric Entropies

### v1

@@ 5 Quantum corrections
We have mainly discussed the canonical bit threads in classical settings. In this section we will include the quantum corrections, then use the positivity of the relative entropy to obtain an energy condition which imposes some constraints on the quantum part. First we will briefly review the relative entropy and the quantum bit threads.
@@
In gravitational settings relative entropy is defined as [35],
@@
where, $H_{\rm{bulk}}$ is the bulk modular Hamiltonian or the gravitational energy, $\lambda$ denotes a one parameter family. Then the positivity of relative entropy tells that
@@
which in gravitational terms can be written as,
@@
Again, by using the CPS formalism, one can rewrite the gravitational energy and entropy as [35],
@@
Therefore the positivity of relative entropy (5) can be expressed in terms of $\mathbf{k}$ as,
@@
The boundary of the region $\Sigma_{A}$ is chosen to be: $\partial\Sigma_{A}=A-m(A)$. This relation will give us the energy condition by incorporating the idea of quantum bit threads.
@@
Quantum bit threads has been introduced recently in [11], and independently in [12]. Although both of them are closely related and can be shown to be the same in the leading orders, here the former one is discussed briefly. The quantum bit thread vector field $v$ is defined as $v=v_{0}+v_{q}$, such that:
@@
Here, $S_{\rm{HEE}}^{0}$ is the classical part of the HEE given by the area of the RT surface and $S_{\rm{bulk}}$ denotes the bulk entanglement entropy arises from quantum corrections. One can separate out the quantum part of $v$ as,
@@
with $\partial\Sigma_{A}=A-m(A)$. The integrand $s_{q}(x)$ was termed as entanglement contour in [11] but the authors emphasised that it need not to be a positive definite quantity as it was originally introduced in [36]. Later we will see that this quantity should be non-negative at least perturbatively in first order. So, to avoid any confusion, we can simply call the quantity $s_{q}(x)$ the ‘bulk entanglement density’ which gives rise to the bulk entanglement. The bulk entropy can also be written as the expectation value of the bulk modular Hamiltonian,
@@
From the above definition it is clear that the divergenceless condition should be modified to incorporate the quantum part of the bit threads. Comparing the above two equations the modified divergence relation is given by,
@@
We will use this modified divergence relation along with the positivity of relative entropy to obtain an energy condition.
@@
The canonical bit threads prescription can be extended to include quantum corrections as well. First let us look at the classical case with matter. Recall the fundamental identity of the CPS formalism,
@@
where, $\delta\boldsymbol{C}_{a}$ is related to the constraints of the theory. In presence of matter, $\delta\boldsymbol{C}_{a}=-\delta T_{ab}\varepsilon^{b}$. So even if the symplectic current vanishes (in the presence of a Killing vector $\xi$), $\mathbf{k}_{\xi}$ does not remain a closed form: $\mathrm{d}\mathbf{k}_{\xi}=-\xi^{a}\delta\boldsymbol{C}_{a}$. The vector $v^{a}_{th}$ constructed from this codim-2 form will also not be divergenceless101010 It implies that similar to the higher curvature gravity case [10], the divergence relation of bit threads should be modified in the presence of classical matter as well.. In the semiclassical regime, we consider the matter to be quantum fields. Then the constraint becomes $\delta\boldsymbol{C}_{a}=-\langle\delta T_{ab}\rangle\varepsilon^{b}$. For exact symmetries, the identity (5.8) reduces to,
@@
This correctly captures the quantum corrections in this regime. Therefore the positivity inequality (5) of relative entropy can be rewritten in terms of bit threads vector field as follows
@@
This relation is true for any arbitrary size of the boundary region $A$. So the above inequality is true for the integrand itself which gives us the energy condition
@@
We find that this energy condition is equivalent to the dominant energy condition (DEC)111111 To check that we use the physical form of the DEC for perturbative matter, $\delta T_{ab}\xi^{a}X^{b}\;\geq 0\,$, with $F^{a}=-\delta T^{a}_{\;\;b}\xi^{b}$ being a future pointing causal vector, and $\xi^{a}X_{a}<0$. For our case let $X^{a}=n^{a}$ is an unit normal timelike vector. Then it satisfies the condition $\xi^{a}n_{a}<0$. Also, as $F^{a}$ and $n^{a}$ both are future pointing and causal, they satisfy $F^{a}n_{a}\leq 0\qquad\Rightarrow\;-\delta T^{a}_{\;\;b}\xi^{b}n_{a}\leq 0\qquad\Rightarrow~~\delta T_{ab}\xi^{b}n^{a}\geq 0\,$, which is nothing but the classical version of the energy condition we have obtained. . So we conclude that if the perturbative bulk stress-energy tensor satisfies the DEC, then the bulk entanglement density should be positive. It subsequently implies the positivity of the relative entropy.
@@ 6 Bit threads in AdS-Vaidya
Till now we have been looking into static spacetimes except the spherical-Rindler case. But we took the near horizon limit (static in nature) which helped us to define a timelike Killing vector and use the covariant phase space formalism to obtain a divergenceless vector field. Then it was shown that this vector field can be thought as a bit thread configuration which gives us the differential entropy of a hole present in the spacetime. Now, we will apply the similar method to certain dynamical spacetimes.
Consider those dynamical spacetimes which can be mapped to static spacetimes via some conformal transformations. It has been showed that for such cases, the conformal relation plays an important role to derive the thermodynamics of the dynamical black hole [37]. Another example of this kind of geometry is the AdS-Vaidya spcaetime. Here, we will look into the AdS-Vaidya spacetime in 3 dimensions. The metric of this spacetime is
@@
Here we consider the mass $M$ to be used for parametric variations. The conformal Killing vector here is: $\xi=(\partial_{w},0,0)$. Using (2.20), we find
@@
where $\delta_{DD}(.)$ represents the Dirac delta function. Here also we can find a perturbative flow by taking the Hodge dual of this form and contracting with the hypersurface metric. There exists a $\boldsymbol{\mathit{j}}_{\xi}$ such that $\mathbf{k}_{\xi}=\pm\hat{\delta}\boldsymbol{\mathit{j}}_{\xi}$ given by,
@@
Now to find the vector $v^{a}_{cps}$, we need to specify a Cauchy surface. On the other hand, as the spacetime is dynamic, the extremal surface need not lie on a Cauchy surface. But to apply our method, we will restrict ourselves to a constant time slice and look into the spacelike extremal surface. We will take an approach similar to [38]. First, we rewrite the metric (6.1) in terms of $t$ and $r$ coordinates:
@@
$\mathit{j}^{wr}_{\xi}=\dfrac{1}{16\pi G_{N}}\dfrac{\mathfrak{C}}{r},$ (6.5)
where $\mathfrak{C}$ is an overall constant which can be fixed by imposing the boundary conditions. The above expression is quite similar to the pure AdS case which tells that the codim-2 current we obtained is correct. Using the map (2.13), we can find a divergenceless vector field,
@@
Now, comparing the vector (6.6) with the above unit normal of RT surface (6.7), one can easily check that $v^{a}_{cps}|_{RT}\neq\mathrm{C}n^{a}$. So, we do need to add $v^{a}_{\mathrm{d}}$ to make the vector field a bit thread flow. We solve the differential equation $\nabla_{a}v^{a}_{\mathrm{d}}=\nabla_{a}\nabla^{a}\varphi=0$ in the background (6.4) with proper boundary conditions, and find
@@
It reduces to the unit normal (6.8) at the RT surface (6.7), which confirms that the above vector field is a bit thread vector. One can also go to the Poincaré and find a flow, for which we have already shown that the vector field obtained from the CPS is not enough. We need to add the term $v^{a}_{\mathrm{d}}$ so that the total vector field becomes a bit thread configuration.
@@
with AdS radius $L=1$. The vector obtain from this form is given by (3.4),
@@
We need to look whether the vector field becomes proportional to the unit normal at the RT surface. We can find the RT surface equation using the geodesics entirely in the BTZ phase and using the metric (6.4). It is given by
@@
where, $F(\mathsf{k}|\mathsf{m})$ is the incomplete elliptic integral of first kind and $E(\mathsf{k}|\mathsf{m})$ is the incomplete elliptic integral of second kind [^12] The incomplete elliptic integrals of first and second kind are given by[39], $F(\mathsf{k}|\mathsf{m})=\int_{0}^{\sin{\mathsf{k}}}\left[(1-t^{2})(1-\mathsf{m}\,t^{2})\right]^{-1/2}\,;\quad E(\mathsf{k}|\mathsf{m})=\int_{0}^{\sin{\mathsf{k}}}(1-t^{2})^{-1/2}(1-\mathsf{m}\,t^{2})^{1/2}.$ . The unit normal to this surface is given by,
@@
From the above expression and (6.11) it is clear that the resultant vector field $v^{a}_{cps}$ is not proportional to the normal $n^{a}$ at the RT surface (6.12). So, we need to add the ambiguity field $v_{\mathrm{d}}^{a}$ to make it a flow for a valid bit thread configuration. One can find this $v_{\mathrm{d}}^{a}$ by using PDE and the boundary conditions as we did for the pure AdS phase of the AdS-Vaidya geometry.
@@
So we see that although the AdS3-Vaidya spacetime does not have a Killing vector, the presence of a conformal Killing vector can provide us the codim-2 forms and the corresponding vector field. This is surprising in a way that even in certain dynamical spacetimes, we can define (perturbative) bit thread flows using conformal Killing vectors rather than genuine Killing vectors. It extends the bit thread formalism beyond static spacetimes, though finding and interpreting the complete non-perturbative flow in general remains an open question.
7 Conclusions and Discussions
@@
In this work, we have explored the the bit thread formalism and its broader applicability to geometric entropy concepts beyond the conventional HEE. Our study has centred on several key themes: formulating methods based on CPS formalism to obtain bit thread configurations, their connections to Wald and differential entropy, a reformulation of thermodynamic first laws in terms of flow dynamics, and preliminary steps toward accommodating quantum corrections and time dependence.
@@
We attempt to construct bit thread configuration using quantities from the CPS formalism section. We found that although it naturally produces a divergenceless vector field $v_{cps}^{a}$, it generally does not meet the norm bound and saturation conditions required of HEE threads $v^{a}$. A gauge freedom $v_{\mathrm{d}}^{a}$ needed to satisfy the constraints of bit threads can be obtained by solving for harmonic function in $3$ dimensional spacetime. We showed that the bit threads vector field and $v^{a}_{cps}$ are in same equivalence class $v^{a}\sim v_{cps}^{a}$. This is one of the main results of our work. In spacetimes where the RT surface becomes horizon, the vector $v^{a}_{cps}$ is same as the bit thread. Then by applying suitable coordinate transformations to it results in a bit thread configuration in the original metric. We used this method to obtain a bit thread configuration for Poincaré AdS in section 2. This method does not require any prior knowledge of the RT surface and its unit normal, but surprisingly we find the resultant thread configuration to be same as the geodesic bit threads [18] up to an overall constant. These alternative ways for bit thread construction may be more suitable for certain setups.
@@
We also extended this flow prescription to rewrite other entropies in the bit thread language. First we related the horizon entropy to this flow picture. In section 3, we showed how bit thread like vector can be found for AdS black holes which gives the Wald entropy once integrated over the bifurcation surface. Same can be done for non-AdS black holes also, as the Wald entropy is defined for any stationary black hole with bifurcation point. Similarly, the cosmological horizon yields a thread description. Another entropy which can be expressed in terms of threads is the differential entropy.
@@
The CPS approach also inspired a new way to view thermodynamic first laws (Section 4). By utilising the Hodge dual of the current 2-form $\mathbf{k}_{\xi}$, we introduced the thermodynamic variation flow $v_{th}^{a}=g^{ab}(\star\mathbf{k}_{\xi})_{b}$, which captures the flux balances associated with entropy changes. These balance equations provide a geometric visualisation of how entropy variations arise from the bulk symplectic current, which is relevant for both black hole (and cosmological horizon) and entanglement first laws. For differential entropy, we find the first law relates the Lorentzian and Riemannian (perturbative) bit threads.
@@
  • Higher curvature gravity: The bit threads picture has been explored in higher curvature gravities in general [9], but an explicit expression for some particular theory is still missing. The method used here to construct a flow using the CPS formalism can be extended to these higher derivative gravities very naturally as the CPS formalism is applicable to more general formulations of gravity [15]. In [9], the authors showed how the norm bound changes for higher curvature gravity, while the vector field remains divergenceless. This property comes naturally following the CPS formalism and one should be able to construct a bit thread configuration by imposing the modified norm bound. It can be further extended to the cubic gravitational theories to possibly address the issue of ‘splitting’ as discussed in [40].
  • Towards covariant bit threads: The notion of covariant bit threads has been explored in [4]. Their exploration is mainly focused on the definition and general properties of the covariant bit threads. Unlike the Riemmanian ones, the covariant threads live in the whole spacetime and its flux is maximum at the HRT surface of the corresponding boundary region. The paper showed that we can find a covariant flow from a closed (d1)(d-1)-form obeying the required norm bound. If one wants to obtain these threads from CPS formalism, the Noether current (B.6) is a promising candidate which is a (d1)(d-1)-form residing in the spacetime unlike the codim-2 current discussed in this paper. Conservation ensures that the Noether current is a closed form (onshell). So one can find a divergenceless vector field by taking the Hodge dual of the Noether current,
@@
    We hope that this vector can be related to the ‘V-threads’ presented in [4] but the exact relation needs further investigation. Finding these threads explicitly will improve our understanding of the connection between entanglement and geometry. This method may also be helpful to extend the bulk reconstruction in the covariant setting similar to the Riemannian case as done in [17].
@@
  • Quantum energy conditions: One interesting direction would be to derive the quantum focussing conjecture [41] and corresponding quantum null energy conditions [42] using this flow based prescription. It will also help to understand the covariant entropy bound and quantum Bousso bound through the bit threads perspective and may shed new light on the underlying structure of these quantities.
@@
  • Generalised entanglement wedge: The construction of a flow from CPS formalism is also applicable to non-holographic spacetimes as discussed earlier. Recently, there has been a proposal of generalised entanglement wedge (GEW) in general gravitational spacetimes [43, 44], which states that for any gravitational bulk region $\mathtt{b}$, there is a GEW $\mathtt{E(b)}$ ($\supset\mathtt{b}$) on a static Cauchy surface. Then the fine grained entropy of the region $\mathtt{b}$ is given by the generalised entropy of the GEW $S_{gen}(\mathtt{E(b)})$. The authors of [45] extended the bit thread formulation for the GEW proposal in general gravitational spacetimes. The construction of bit threads using the CPS formalism provides a canonical way to extend the flow prescription to this GEW proposal.
@@
  • Timelike entanglement entropy: It would be interesting to find a flow based picture for the timelike entanglement entropy [46]. One possible way is to use Wick rotated Euclidean metric in AdS and apply the method described in section 2. It may also be related to the pseudo entropy [47] which is more general than the timelike EE.
@@
  • Connection to geometric modular flows: In CFTs, entanglement entropy of a subregion can be obtained by using the notion of modular geometric flows [48]. In the holographic context, the same can be obtained by using the bit threads. A possible connection has been already mentioned in [48]. The difference between these two formalism is that the geometric modular flow is uniquely determined whereas the configuration of bit threads of highly non-unique. Apparently, the bit threads and the geometric modular flows are normal to each other at the boundary subregion. If we consider the case in a Rindler wedge, the CPS method provides a canonical thread configuration. It will be worth exploring the connection between these two kind of flows in future.
Acknowledgements

### v2

@@ 5 Quantum corrections
We have mainly discussed the canonical bit threads in classical settings. In this section, we will include the quantum corrections, then use the positivity of the relative entropy to obtain an energy condition that imposes constraints on the quantum part. First, we will briefly review the relative entropy and the quantum bit threads.
@@
In gravitational settings, relative entropy is defined as [41],
@@
where, $H_{\rm{bulk}}$ is the bulk modular Hamiltonian or the gravitational energy, [^13] It is same as the modular energy $E$ discussed in the previous section. $\lambda$ denotes a one-parameter family. Then the positivity of relative entropy tells us that
@@
which, in gravitational terms, is given by
@@
Again, by using the CPS formalism, one can rewrite the above inequality consisting of the gravitational energy and entropy as [41],
@@
Therefore, the positivity of relative entropy (5) can be expressed in terms of $\mathbf{k}$ as follows
@@
The boundary of the region $\Sigma_{A}$ is chosen to be: $\partial\Sigma_{A}=A-m(A)$. The above relation will provide us with an energy condition, along with incorporating the concept of quantum bit threads.
@@
Quantum bit threads have been introduced recently in [4], and independently in [51]. Although both are closely related and can be shown to be the same in leading order, here we will briefly discuss the former one. The quantum bit thread vector field $v$ is defined as $v=v_{0}+v_{q}$, such that:
@@
Here, $S_{\rm{HEE}}^{0}$ is the classical part of the HEE given by the area of the RT surface, and $S_{\rm{bulk}}$ denotes the bulk entanglement entropy arising from quantum corrections. One can separate the quantum part of $v$ as,
@@
with $\partial\Sigma_{A}=A-m(A)$. The integrand $s_{q}(x)$ was termed the entanglement contour in [4], but the authors emphasised that it need not be a positive definite quantity as it was originally introduced in [55]. However, we find that it is non-negative perturbatively in the first order. So, to avoid any confusion, we will call the quantity $s_{q}(x)$ the ‘bulk entanglement density’, which gives rise to the bulk entanglement. The bulk entropy can also be written as the expectation value of the bulk modular Hamiltonian,
@@
Here, $n_{a}\wedge\varepsilon^{a}={\boldsymbol{\epsilon}}$, where $\boldsymbol{\epsilon}$} is the volume form for $d$-dimensional spacetime. From the above definition, it is clear that the divergenceless condition should be modified to incorporate the quantum part of the bit threads, and the modified divergence relation is given by,
@@
The canonical bit threads prescription can be extended to include quantum corrections as well. First, let us examine the classical case with matter. Recall the fundamental identity of the CPS formalism,
@@
where, $\delta\boldsymbol{C}_{a}$ is related to the constraints of the theory. In presence of matter, $\delta\boldsymbol{C}_{a}=-\delta T_{ab}\varepsilon^{b}$. So even if the symplectic current vanishes (in the presence of a Killing vector $\xi$), $\mathbf{k}_{\xi}$ does not remain a closed form: $\mathrm{d}\mathbf{k}_{\xi}=-\xi^{a}\delta\boldsymbol{C}_{a}$. The vector $v^{a}_{th}$ constructed from this codim-2 form will also not be divergenceless. [^14] It implies that, similar to the higher curvature gravity case [31], the divergence relation of bit threads should be modified in the presence of classical matter as well. In the semiclassical regime, we consider the matter to be composed of quantum fields. Then the constraint becomes $\delta\boldsymbol{C}_{a}=-\langle\delta T_{ab}\rangle\varepsilon^{b}$. For exact symmetries, the identity (5.8) reduces to,
@@
This correctly captures the quantum corrections in this regime. Therefore, the positivity inequality (5) of relative entropy can be rewritten in terms of the bit threads vector field as follows
@@
This relation is true for any arbitrary size of the boundary region $A$. So the above inequality is true for the integrand itself, which gives us the energy condition
@@
We find that this energy condition is equivalent to the dominant energy condition (DEC). [^15] To check it, we use the physical form of the DEC for perturbative matter $\delta T_{ab}\xi^{a}X^{b}\;\geq 0\,$, with $F^{a}=-\delta T^{a}_{\;\;b}\xi^{b}$ being a future pointing causal vector, and $\xi^{a}X_{a}<0$. For our case, let $X^{a}=n^{a}$ be a unit normal timelike vector. Then it satisfies the condition $\xi^{a}n_{a}<0$. Also, as $F^{a}$ and $n^{a}$ both are future-pointing and causal, they satisfy $F^{a}n_{a}\leq 0\qquad\Rightarrow\;-\delta T^{a}_{\;\;b}\xi^{b}n_{a}\leq 0\qquad\Rightarrow~~\delta T_{ab}\xi^{b}n^{a}\geq 0\,$, which is nothing but the classical version of the energy condition that we have obtained. So we conclude that if the perturbative bulk stress-energy tensor satisfies the DEC, then the bulk entanglement density should be positive: $\delta s_{q}(x)\geq 0$. It subsequently implies the positivity of the relative entropy.
@@ 6 Bit threads in AdS-Vaidya
In this section, we attempt to apply a similar method to the one employed in the previous sections to certain dynamical spacetimes.
Consider those dynamical spacetimes that can be mapped to static spacetimes via some conformal transformations. It has been shown that for such cases, the conformal relation plays an important role in deriving the thermodynamics of the dynamical black holes [47]. Another example of this kind of geometry is the AdS-Vaidya spacetime. Here, we will look into the AdS-Vaidya spacetime in 3 dimensions. The metric of this spacetime is
@@
Here, we consider the mass $M$ to be used for parametric variations. The conformal Killing vector here is: $\xi=(\partial_{w},0,0)$. Using (2.20), we find
@@
where $\delta_{DD}(.)$ represents the Dirac delta function. We can find a perturbative flow by taking the Hodge dual of this form and contracting with the hypersurface metric. There exists a $\boldsymbol{\mathit{j}}_{\xi}$ such that $\mathbf{k}_{\xi}=\pm\hat{\delta}\boldsymbol{\mathit{j}}_{\xi}$ given by,
@@
Now, to find the vector $v^{a}_{cps}$, we need to specify a Cauchy surface. On the other hand, as the spacetime is dynamic, the extremal surface need not lie on a Cauchy surface. However, to apply our method, we will restrict ourselves to a constant time slice and examine the spacelike extremal surfaces. We will adopt an approach similar to that of [15]. First, we rewrite the metric (6.1) in terms of $t$ and $r$ coordinates:
@@
$\mathit{j}^{tr}_{\xi}=\dfrac{1}{16\pi G_{N}}\dfrac{\mathfrak{C}}{r},$ (6.5)
where $\mathfrak{C}$ is an overall constant that can be fixed by imposing the boundary conditions. The above expression is quite similar to the pure AdS case, which suggests that the codim-2 current we have obtained is correct. Using the map (2.13), we can find a divergenceless vector field,
@@
Now, comparing the vector (6.6) with the above unit normal of RT surface (6.7), one can easily check that $v^{a}_{cps}|_{RT}\neq\mathrm{C}n^{a}$. So, we need to add $v^{a}_{\mathrm{d}}$ to make the vector field a bit thread flow. We solve the differential equation $\nabla_{a}v^{a}_{\mathrm{d}}=\nabla_{a}\nabla^{a}\varphi=0$ in the background (6.4) with proper boundary conditions, and find
@@
It reduces to the unit normal (6.8) at the RT surface (6.7), which confirms that the above vector field is a bit thread vector field. One can also go to the Poincaré coordinates and find a flow, for which we have already shown that the vector field obtained from the CPS is not enough. We need to add the term $v^{a}_{\mathrm{d}}$ so that the total vector field becomes a bit thread configuration.
@@
with AdS radius $L=1$. The vector obtained from this form is given by (3.4),
@@
We need to look at whether the vector field becomes proportional to the unit normal at the RT surface. We can find the RT surface equation using the geodesics entirely in the BTZ phase and using the metric (6.4). It is given by
@@
where $F(\mathsf{k}|\mathsf{m})$ is the incomplete elliptic integral of the first kind and $E(\mathsf{k}|\mathsf{m})$ is the incomplete elliptic integral of the second kind. [^16] The incomplete elliptic integrals of the first and second kind are given by [1], $F(\mathsf{k}|\mathsf{m})=\int_{0}^{\sin{\mathsf{k}}}\left[(1-t^{2})(1-\mathsf{m}\,t^{2})\right]^{-1/2}\,;\quad E(\mathsf{k}|\mathsf{m})=\int_{0}^{\sin{\mathsf{k}}}(1-t^{2})^{-1/2}(1-\mathsf{m}\,t^{2})^{1/2}.$ The unit normal to this surface is given by,
@@
From the above expression and (6.11), it is clear that the resultant vector field $v^{a}_{cps}$ is not proportional to the normal $n^{a}$ at the RT surface (6.12). So, we need to add the ambiguity field $v_{\mathrm{d}}^{a}$ to make it a flow for a valid bit thread configuration. One can find this $v_{\mathrm{d}}^{a}$ by using the method of characteristics with proper boundary conditions, as we did for the pure AdS phase of the AdS-Vaidya geometry.
@@
So we see that although the AdS3-Vaidya spacetime does not have a Killing vector, the presence of a conformal Killing vector can provide us with the codim-2 forms and the corresponding vector field. This is surprising in a way that even in certain dynamical spacetimes, we can define (perturbative) bit thread flows using conformal Killing vectors rather than genuine Killing vectors. It extends the bit thread formalism beyond static spacetimes, though finding and interpreting the complete non-perturbative flow in general remains an open question.
7 Conclusion and Discussion
@@
In this work, we have explored the bit thread formalism and its broader applicability to geometric entropy concepts beyond the conventional HEE. Our study has centred on several key themes: formulating methods based on CPS formalism to obtain bit thread configurations, their connections to Wald and differential entropy, a reformulation of thermodynamic first laws in terms of flow dynamics, and preliminary steps toward accommodating quantum corrections and time dependence.
@@
We attempt to construct a bit thread configuration using quantities from the CPS formalism. We found that although it naturally produces a divergenceless vector field $v_{cps}^{a}$, it generally does not meet the norm bound and saturation conditions required of HEE threads $v^{a}$. A gauge freedom $v_{\mathrm{d}}^{a}$ needed to satisfy the constraints of bit threads can be obtained by solving for a harmonic function in $3$-dimensional spacetime. We showed that the bit threads vector field and $v^{a}_{cps}$ are in the same equivalence class $v^{a}\sim v_{cps}^{a}$. This is one of the main results of our work. In spacetimes where the RT surface becomes a horizon, the vector $v^{a}_{cps}$ is same as the bit thread vector field. Then, applying suitable coordinate transformations to it results in a bit thread configuration in the original metric. We used this method to obtain a bit thread configuration for Poincaré AdS in section 2. This method does not require any prior knowledge of the RT surface and its unit normal; however, surprisingly, we find that the resultant thread configuration is same as the geodesic bit threads [3] up to an overall constant. These alternative ways for bit thread construction may be more suitable for specific setups.
@@
We have also extended this flow prescription to rewrite other entropies in the bit thread language. First, we related the horizon entropy to this flow picture. In section 3, we showed how a bit thread-like vector can be found for AdS black holes, which gives the Wald entropy once integrated over the bifurcation surface. Another entropy that can be expressed in terms of threads is the differential entropy for certain spacetimes, which is also defined for non-holographic spacetimes.
@@
The CPS approach also inspired a new way to view thermodynamic first laws (Section 4). By utilising the Hodge dual of the current codim-2 form $\mathbf{k}_{\xi}$, we introduced the thermodynamic variation flow $v_{th}^{a}=g^{ab}(\star\mathbf{k}_{\xi})_{b}$, which captures the flux balances associated with entropy changes. These balance equations provide a geometric visualisation of how entropy variations arise from the bulk symplectic current, which is relevant for both black hole thermodynamics and the entanglement entropy first law. For differential entropy, we find that the first law relates the Lorentzian and Riemannian (perturbative) bit threads.
@@
  • Higher curvature gravity: The bit threads picture has been explored in higher curvature gravities in general [31], but an explicit construction for some particular theory is still missing. The method used here to construct a flow using the CPS formalism can be extended to these higher-derivative gravities very naturally, as the CPS formalism is applicable to more general formulations of gravity [39]. In [31], the authors showed how the norm bound changes for higher curvature gravity, while the vector field remains divergenceless. This property comes naturally following the CPS formalism, and one should be able to construct a bit thread configuration by imposing the modified norm bound. It can be further extended to the cubic gravitational theories to possibly address the issue of ‘splitting’ as discussed in [14].
  • Towards covariant bit threads: The notion of covariant bit threads has been explored in [34]. Their exploration is mainly focused on the definition and general properties of the covariant bit threads. Unlike the Riemannian ones, the covariant threads live in the whole spacetime, and their flux is maximum at the HRT surface of the corresponding boundary region. The paper showed that we can find a covariant flow from a closed (d1)(d-1)-form obeying the required norm bound. If one wants to obtain these threads from CPS formalism, the Noether current (B.6) is a promising candidate, which is a (d1)(d-1)-form residing in the spacetime. Conservation ensures that the Noether current is a closed form (on-shell). So one can find a divergenceless vector field by taking the Hodge dual of the Noether current,
@@
    We hope that this vector field can be related to the ‘V-threads’ presented in [34], but the exact relation needs further investigation. Identifying these threads explicitly will enhance our understanding of the relationship between entanglement and geometry. This method may also be helpful to extend the bulk reconstruction in the covariant setting, similar to the Riemannian case, as done in [2].
@@
  • Quantum energy conditions: One interesting direction would be to derive the quantum focusing conjecture [11] and corresponding quantum null energy conditions [10] using this flow-based prescription. It will also help to understand the covariant entropy bound and quantum Bousso bound through the bit threads perspective, and may shed new light on the underlying structure of these quantities.
@@
  • Generalised entanglement wedge: The construction of a flow from CPS formalism is also applicable to general gravitational spacetimes. Recently, there has been a proposal of generalised entanglement wedge (GEW) in general gravitational spacetimes [12, 13], which states that for any gravitational bulk region $\mathtt{b}$, there is a GEW $\mathtt{E(b)}$ ($\supset\mathtt{b}$) on a static Cauchy surface. Then the fine-grained entropy of the region $\mathtt{b}$ is given by the generalised entropy of the GEW $S_{gen}(\mathtt{E(b)})$. The authors of [24] extended the bit thread formulation for the GEW proposal in general gravitational spacetimes. The construction of bit threads using the CPS formalism provides a canonical way to extend the flow prescription to this GEW proposal.
@@
  • Timelike entanglement entropy: It would be interesting to find a flow-based picture for the timelike entanglement entropy [22]. One possible approach is to use the Wick-rotated Euclidean metric in AdS and apply the method described in section 2. It may also be related to the pseudo entropy [46], which is more general than the timelike EE.
@@
  • Connection to geometric modular flows: In CFTs, entanglement entropy of a subregion can be obtained by using the notion of modular geometric flows [16]. In the holographic context, the same can be obtained by using the bit threads. A possible connection has already been mentioned in [16]. The difference between these two formalisms is that the geometric modular flow is uniquely determined, whereas the configuration of bit threads is highly non-unique. Apparently, the bit threads and the geometric modular flows are normal to each other at the boundary subregion. If we consider the case in the AdS-Rindler wedge, the CPS method provides a canonical thread configuration. It will be worth exploring the connection between these two kinds of flows in future.
Acknowledgments